# Model-assisted inference for dynamic causal effects in staggered rollout cluster randomized experiments

### v2

@@ 5.1 Simulation study I
We compare nine estimators, using individual-level data: $\widehat{\tau}_{j,\mathrm{I}}$, $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}$, $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}[C]$, and $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{ancova}}$; using cluster-period averages: $\widehat{\tau}_{j,\mathrm{A}}$ and $\widehat{\tau}_{j,\mathrm{A}}^{\mathrm{adj}}$; and using scaled cluster-period totals: $\widehat{\tau}_{j,\mathrm{T}}$, $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}[\pi]$, and $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}[\pi,\pi C]$. For estimators with covariate adjustment, $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}$ and $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{ancova}}$ adjust for the centered covariate $X_{ijk}^{c}$, $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}[C]$ and $\widehat{\tau}_{j,\mathrm{A}}^{\mathrm{adj}}$ adjust for the cluster-level centered covariate $\overline{X}_{ij}^{c}$, $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}[\pi]$ adjusts for the centered version of the normalized cluster-period weight $\pi_{ij\cdot}=N_{ij}/N_{j}$, and $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}[\pi,\pi C]$ adjusts for the centered version of $(\pi_{ij\cdot},\pi_{ij\cdot}C_{ij})^{T}$, where $C_{ij}=I\overline{X}_{ij}$. Under this setting, $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}[\pi,\pi C]$ is equivalent to $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}[\pi,\widetilde{X}$] which adjusts for the centered $(\pi_{ij\cdot},\widetilde{X}_{ij})^{T}$ where $\widetilde{X}_{ij}=I\pi_{ij\cdot}\overline{X}_{ij}=\pi_{ij\cdot}C_{ij}$.
@@ 5.3 Additional simulation studies
Additional simulations are to further illustrate properties of the proposed estimators. In simulation study III, we give a counter-example that covariate adjustment does not necessarily improve the performance of $\widehat{\tau}_{j,\mathrm{I}}$ and $\widehat{\tau}_{j,\mathrm{A}}$. In simulation study IV, we investigate the finite-sample performance of the estimators with a large but not dominant cluster. Results from simulation studies III and IV are presented in Sections S6 and S7 of the Online Supplement. Complementary simulations are carried out under the same data-generating processes, but with $I=60$ to examine finite-sample performance. The results of these simulations are similar and given in Section S7 of the Online Supplement.
@@ 6 Data example
We illustrate the proposed methods by analyzing the Acute Coronary Syndrome Quality Improvement in Kerala (ACS QUIK) trial data (Huffman et al., 2018). The ACS QUIK trial is a cross-sectional SR-CRE conducted to evaluate the effect of a quality improvement tool kit on health outcomes in patients with acute myocardial infarction. The cluster-level treatment was rolled out to participating hospitals; in the first four-month period, no hospitals were given the tool kit; in the last four-month period, all hospitals received the tool kit, and the entire rollout consisted of the four four-month periods in between.
@@
We consider two sets of individual weights, i.e., $w_{ijk}=1$ and $w_{ijk}=N_{ij}^{-1}$. The former leads to the DWATE estimand in period $j$ defined as the pooled average, $\tau_{j}(a,a^{\prime})=N_{j}^{-1}\sum_{i=1}^{I}\sum_{k=1}^{N_{ij}}\{Y_{ijk}(a)-Y_{ijk}(a^{\prime})\}$, and the latter accounts for the imbalance in cluster sizes and defined an average of cluster means, $\tau_{j}(a,a^{\prime})=I^{-1}\sum_{i=1}^{I}\{\overline{Y}_{ij\cdot}(a)-\overline{Y}_{ij\cdot}(a^{\prime})\}$ with $\overline{Y}_{ij\cdot}(a)=N_{ij}^{-1}\sum_{i=1}^{I}Y_{ijk}(a)$ and $\overline{Y}_{ij\cdot}(a^{\prime})=N_{ij}^{-1}\sum_{i=1}^{I}Y_{ijk}(a^{\prime})$. We estimate these estimands using the seven methods developed previously, using individual data: $\widehat{\tau}_{j,\mathrm{I}}$, $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}$, and $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{ancova}}$; using cluster-period averages: $\widehat{\tau}_{j,\mathrm{A}}^{\mathrm{adj}}$; and using scaled cluster-period totals: $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}$, $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}[\pi]$, and $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}[\pi,\pi C]$, with results presented in Table 4. Under the weight $w_{ijk}=1$, only $\widehat{\tau}_{j,\mathrm{A}}^{\mathrm{adj}}$ returned with an estimate of the $\mathrm{OWTE}^{sim}$ that is significantly greater than zero, while all seven estimators yielded statistically significant estimates of the $\mathrm{OAWTE}^{sim}$. Under the weight $w_{ijk}=N_{ij}^{-1}$, no estimators yielded statistically significant estimates of $\mathrm{OWTE}^{sim}$ or $\mathrm{OAWTE}^{sim}$. However, under the weight $w_{ijk}=1$, because of the large variation in cluster-period sizes (coefficients of variation of cluster-period sizes are near one for all four periods), the regularity condition (C2) in Section S2 of the Online Supplement requiring no clusters to have weights significantly dominant over others in each period is likely violated. This could result in the asymptotic normality of the regression estimators and the conservativeness of variance estimators both being compromised, as demonstrated by the simulation results in Section S7 of the Online Supplement when a large cluster (cluster weight imbalance) exists and the simulation results in Su and Ding (2021, §6.5) when a dominant cluster is present in a parallel-arm CRE. This situation can be avoided under the weight $w_{ijk}=N_{ij}^{-1}$ because it gives identical cluster weights, satisfying the aforementioned regularity condition (C2) and leading to more trustworthy variance estimates. Therefore, combining these results, it could be concluded that the toolkit intervention did not have a significant treatment or placebo effect in reducing the risk of 30-day major adverse cardiovascular events during the rollout.
@@ 7 Concluding remarks
We provide a comprehensive treatment of causal inference for the class of staggered rollout cluster randomized experiments, including a class of causal estimands accounting for adoption and calendar times, as well as regression estimators using different levels of data. To pursue our estimands, in each calendar period, we can first compare the difference between any pair of treatment adoption times by estimating the DWATE estimand. Other collections of estimands, including weighted average treatment effects and anticipated weighted average treatment effects (and examples listed in Section S1 of the Online Supplement), are linear combinations of the DWATEs. Regression estimators using individual-level data, cluster-period averages, and scaled cluster-period totals are formally characterized, which, under the design-based perspective, are model-assisted. We established the consistency and asymptotic normality for each regression estimator and proved that the CR and HC variance estimators are asymptotically conservative. Relatedly, our results on the conservativeness of the HC variance estimator also provide a rigorous theoretical foundation for its adoption in design-based analyses of SR-IREs (Roth and Sant’Anna, 2023). Through analytic comparisons, we found that the regression estimators using individual-level data and cluster-period averages can be suboptimal, whereas the estimator using scaled cluster-period totals and adjusting for covariates is more efficient with a smaller asymptotic variance. This implication, in fact, also offers novel insights into the analysis of cluster randomized DID; that is, if a cluster randomized DID design only has access to cluster-period level data, one should opt for collecting cluster-period totals rather than cluster-period averages, such that the analysis is asymptotically more efficient. Our findings generalize those previously developed for much simpler parallel-arm CREs in Su and Ding (2021) to a more complex class of SR-CREs. This generalization, however, is not trivial because our asymptotic results are established under the randomization distribution induced by assigning clusters to different adoption times and jointly for a finite-dimensional vector of DWATE estimands. The conservativeness of the CR and HC variance estimators, hence, is in the sense of Löwner ordering and holds for inference about any linear combinations of the DWATE estimands.
@@
We acknowledge that our asymptotic theory assumes the number of clusters to approach infinity with bounded cluster-period sizes and a fixed number of calendar periods. This asymptotic regime is standard and follows the ones used in Middleton and Aronow (2015); Su and Ding (2021); Chen and Li (2024). Nevertheless, with SR-CREs, this is not the only possible asymptotic regime, and it will be interesting to develop complementary asymptotic theories under the alternative regime where the number of clusters stays fixed, but the cluster size approaches infinity or both approach infinity, reminiscent of those studied in Xie and Yang (2003) for generalized estimating equation estimators.

### v3

@@ 5.1 Simulation study I
We compare nine estimators, using individual-level data: $\widehat{\tau}_{j,\mathrm{I}}$, $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}$, $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}[C]$, and $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{ancova}}$; using cluster-period averages: $\widehat{\tau}_{j,\mathrm{A}}$ and $\widehat{\tau}_{j,\mathrm{A}}^{\mathrm{adj}}$; and using scaled cluster-period totals: $\widehat{\tau}_{j,\mathrm{T}}$, $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}[\pi]$, and $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}[\pi,\pi C]$. For estimators with covariate adjustment, $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}$ and $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{ancova}}$ adjust for the centered covariate $X_{ijk}^{c}$, $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}[C]$ and $\widehat{\tau}_{j,\mathrm{A}}^{\mathrm{adj}}$ adjust for the cluster-level centered covariate $\overline{X}_{ij}^{c}$, $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}[\pi]$ adjusts for the centered version of the normalized cluster-period weight $\pi_{ij\cdot}=N_{ij}/N_{j}$, and $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}[\pi,\pi C]$ adjusts for the centered version of $(\pi_{ij\cdot},\pi_{ij\cdot}C_{ij})^{\top}$, where $C_{ij}=I\overline{X}_{ij}$. Under this setting, $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}[\pi,\pi C]$ is equivalent to $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}[\pi,\widetilde{X}$] which adjusts for the centered $(\pi_{ij\cdot},\widetilde{X}_{ij})^{\top}$ where $\widetilde{X}_{ij}=I\pi_{ij\cdot}\overline{X}_{ij}=\pi_{ij\cdot}C_{ij}$.
@@ 5.3 Additional simulation studies
Additional simulations are to further illustrate properties of the proposed estimators. In simulation study III, we provide a counterexample showing that covariate adjustment does not necessarily improve the performance of $\widehat{\tau}_{j,\mathrm{I}}$ and $\widehat{\tau}_{j,\mathrm{A}}$. In simulation study IV, we investigate the finite-sample performance of the estimators with a large but not dominant cluster. Results from simulation studies III and IV are presented in Sections S7 and S8 of the Online Supplement. Complementary simulations are carried out under the same data-generating processes, but with $I=60$ to examine finite-sample performance. The results of these simulations are presented in Section S8 of the Online Supplement and are similar.
@@ 6 Data example
We illustrate the proposed methods by analyzing the Acute Coronary Syndrome Quality Improvement in Kerala (ACS QUIK) trial data (Huffman et al., 2018). The ACS QUIK trial is a cross-sectional SR-CRE conducted to evaluate the effect of a quality improvement toolkit (the intervention) on health outcomes in patients with acute myocardial infarction. The cluster-level treatment was rolled out to participating hospitals; in the first four-month period, no hospitals received the toolkit; in the last four-month period, all hospitals received the toolkit, and the entire rollout consisted of the four four-month periods in between.
@@
We consider two sets of individual weights, i.e., $w_{ijk}=1$ and $w_{ijk}=N_{ij}^{-1}$. The former leads to the DWATE estimand in period $j$ defined as the pooled average, $\tau_{j}(a,a^{\prime})=N_{j}^{-1}\sum_{i=1}^{I}\sum_{k=1}^{N_{ij}}\{Y_{ijk}(a)-Y_{ijk}(a^{\prime})\}$, and the latter accounts for the imbalance in cluster sizes and defined an average of cluster means, $\tau_{j}(a,a^{\prime})=I^{-1}\sum_{i=1}^{I}\{\overline{Y}_{ij\cdot}(a)-\overline{Y}_{ij\cdot}(a^{\prime})\}$ with $\overline{Y}_{ij\cdot}(a)=N_{ij}^{-1}\sum_{i=1}^{I}Y_{ijk}(a)$ and $\overline{Y}_{ij\cdot}(a^{\prime})=N_{ij}^{-1}\sum_{i=1}^{I}Y_{ijk}(a^{\prime})$. We estimate these estimands using the seven methods developed previously, using individual data: $\widehat{\tau}_{j,\mathrm{I}}$, $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}$, and $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{ancova}}$; using cluster-period averages: $\widehat{\tau}_{j,\mathrm{A}}^{\mathrm{adj}}$; and using scaled cluster-period totals: $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}$, $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}[\pi]$, and $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}[\pi,\pi C]$, with results presented in Table 4. Under the weight $w_{ijk}=1$, only $\widehat{\tau}_{j,\mathrm{A}}^{\mathrm{adj}}$ returned with an estimate of the $\mathrm{OWTE}^{sim}$ that is significantly greater than zero, while six out of seven estimators yielded statistically significant estimates of the $\mathrm{OAWTE}^{sim}$. Under the weight $w_{ijk}=N_{ij}^{-1}$, no estimators yielded statistically significant estimates of $\mathrm{OWTE}^{sim}$ or $\mathrm{OAWTE}^{sim}$. However, under the weight $w_{ijk}=1$, because of the large variation in cluster-period sizes (coefficients of variation of cluster-period sizes are near one for all four periods), the regularity condition (C2) in Section S2 of the Online Supplement, requiring no clusters to have weights significantly dominant over others in each period, is likely violated. This could result in the asymptotic normality of the regression estimators and the conservativeness of variance estimators both being compromised, as demonstrated by the simulation results in Section S8 of the Online Supplement when a large cluster (cluster weight imbalance) exists and the simulation results in Su and Ding (2021, §6.5) when a dominant cluster is present in a parallel-arm CRE. This situation can be avoided under the weight $w_{ijk}=N_{ij}^{-1}$ because it yields identical cluster weights, satisfying the aforementioned regularity condition (C2) and resulting in more trustworthy variance estimates. Therefore, combining these results, it could be concluded that the toolkit intervention did not have a significant treatment or placebo effect in reducing the risk of 30-day major adverse cardiovascular events during the rollout.
@@ 7 Concluding remarks
We provide a comprehensive treatment of causal inference for the class of staggered rollout cluster randomized experiments, including a class of causal estimands accounting for adoption and calendar times, as well as regression estimators using different levels of data. To pursue our estimands, in each calendar period, we can first compare the difference between any pair of treatment adoption times by estimating the DWATE estimand. Other collections of estimands, including weighted average treatment effects and anticipated weighted average treatment effects (and examples listed in Section S1 of the Online Supplement), are linear combinations of the DWATEs. Regression estimators using individual-level data, cluster-period averages, and scaled cluster-period totals are formally characterized, which, under the design-based perspective, are model-assisted. We established the consistency and asymptotic normality for each regression estimator and proved that the CR and HC variance estimators are asymptotically conservative. Relatedly, our results on the conservativeness of the HC variance estimator also provide a rigorous theoretical foundation for its adoption in design-based analyses of SR-IREs (Roth and Sant’Anna, 2023). Through analytical comparisons, we found that regression estimators using individual-level data and cluster-period averages can be suboptimal, whereas the estimator using scaled cluster-period totals and adjusting for covariates is more efficient, with a smaller asymptotic variance. This implication, in fact, also offers novel insights into the analysis of cluster randomized DID. That is, if a cluster randomized DID design only has access to cluster-period level data, one should opt for collecting cluster-period totals rather than cluster-period averages, thereby making the analysis asymptotically more efficient. Our findings generalize those previously developed for much simpler parallel-arm CREs in Su and Ding (2021) to a more complex class of SR-CREs. This generalization, however, is not trivial because our asymptotic results are established under the randomization distribution induced by assigning clusters to different adoption times and jointly for a finite-dimensional vector of DWATE estimands. The conservativeness of the CR and HC variance estimators, hence, is in the sense of Löwner ordering and holds for inference about any linear combinations of the DWATE estimands.
@@
We acknowledge that our asymptotic theory assumes the number of clusters approaches infinity, with bounded cluster-period sizes and a fixed number of calendar periods. This asymptotic regime is standard and follows the ones used in Middleton and Aronow (2015); Su and Ding (2021); Chen and Li (2024). Nevertheless, with SR-CREs, this is not the only possible asymptotic regime, and it will be interesting to develop complementary asymptotic theories under the alternative regime where the number of clusters stays fixed. Still, the cluster size approaches infinity or both approach infinity, reminiscent of those studied in Xie and Yang (2003) for generalized estimating equation estimators.